### v3

# Born a Transformer – Always a Transformer? On the Effect of Pretraining on Architectural Abilities

Abstract
Transformers have theoretical limitations in modeling certain sequence-to-sequence tasks, yet it remains largely unclear if these limitations play a role in large-scale pretrained LLMs, or whether LLMs might effectively overcome these constraints in practice due to the scale of both the models themselves and their pretraining data. We explore how these architectural constraints manifest after pretraining, by studying a family of retrieval and copying tasks inspired by Liu et al. (2024a). We use a recently proposed framework for studying length generalization (Huang et al., 2025) to provide guarantees for each of our settings. Empirically, we observe an induction-versus-anti-induction asymmetry, where pretrained models are better at retrieving tokens to the right (induction) rather than the left (anti-induction) of a query token. This asymmetry disappears upon targeted fine-tuning if length-generalization is guaranteed by theory. Mechanistic analysis reveals that this asymmetry is connected to the differences in the strength of induction versus anti-induction circuits within pretrained transformers. We validate our findings through practical experiments on real-world tasks demonstrating reliability risks. Our results highlight that pretraining selectively enhances certain transformer capabilities, but does not overcome fundamental length-generalization limits [^1] https://github.com/lacoco-lab/always_a_transformer.
1 Introduction
Transformers (Vaswani et al., 2017) have become the backbone of most LLMs (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023). Given the widespread usage of LLMs, there has been a long line of research to understand their theoretical underpinnings. Early analyses showed that transformers can approximate any sequence-to-sequence function at fixed input length (Yun et al., 2019) and even simulate Turing machines (Pérez et al., 2019) when given unbounded compute. However, subsequent studies have uncovered fundamental expressibility and learnability limitations, finding classes of problems which transformers cannot express across arbitrary input lengths (e.g. Hahn, 2020; Strobl et al., 2024; Merrill and Sabharwal, 2023; Sanford et al., 2024), or which transformers systematically struggle to learn and generalize on (e.g. Hahn and Rofin, 2024; Zhou et al., 2024). Empirical work confirms that small transformers trained from scratch do instantiate such limitations (e.g. Bhattamishra et al., 2020; Huang et al., 2025; Delétang et al., 2023; Zhou et al., 2024). For instance, transformers trained from scratch struggle to generalize to longer or more challenging examples even on tasks as simple as copying a long string with repeated symbols (Zhou et al., 2024), or retrieving which token followed the last occurrence of some query token in a long context (Liu et al., 2024a) — findings in close agreement with theoretical arguments about transformer’s learning biases (Huang et al., 2025).
However, it remains unclear if such results on expressiveness and generalization have any bearing on large-scale pretrained models. Modern LLMs have many more layers, and many orders of magnitude more parameters than the small-scale transformers typically used for testing theoretical limitations. Model scale alone might make theoretical limitations irrelevant at realistic input lengths. Perhaps even more importantly, LLMs are pretrained on a massive scale, with trillions of tokens of varied data that may impart fundamentally new inductive biases and algorithmic abilities (Rytting and Wingate, 2021) that may be invisible in models trained from scratch on the target task alone (Amos et al., 2024). Indeed, LLMs’ capabilities scale in ways that are difficult to foresee: empirical “laws” relate performance to model, data, and compute budget (Kaplan et al., 2020; Hoffmann et al., 2022) and qualitatively new behaviors such as in-context learning, chain-of-thought reasoning, and tool use emerge abruptly at scale (Wei et al., 2022; Ganguli et al., 2022; Bubeck et al., 2023). Since pretraining reshapes both the model’s inductive biases and the sub-circuits it relies on, length-generalization failures observed in small models might not transfer automatically to their large-scale counterparts.
Scope. We focus on retrieval and copying as fundamental operations that provide a test bed for answering these questions. These operations are critical building blocks for many real-world LLM applications, such as question answering and summarization (Wiegreffe et al., 2025; Fan et al., 2024). On a mechanistic level, retrieving from context via induction heads is key to performing in-context learning (Olsson et al., 2022; Song et al., 2025; Crosbie and Shutova, 2025), while faithfully copying exact spans makes retrieval-augmented generation more effective (Yu et al., 2024). Understanding the potential impact of architectural limitations on LLMs is thus of great interest for these tasks.
In this work, we revisit the theory of length generalization for these operations, ask how its predictions hold up for pretrained transformers, and diagnose which abilities are amplified by pretraining and which limitations persist. We study a family of retrieval and copying tasks222i.e., copying of the input, not the training data, similar to Jelassi et al. (2024) (Figure 1) inspired by Liu et al. (2024a); Zhou et al. (2024). We organize retrieval into two umbrella categories of unique and non-unique induction, respectively, in which the query token appears exactly once or multiple times respectively. In each variant, the model must return a token located either before or after a designated occurrence of that query token. We make an analogous distinction in copying between copying unique tokens or repeated tokens, and between copying in the forward or the reverse direction.
All retrieval and copying tasks we consider are expressible by transformers in principle; however, theory predicts systematic differences in length generalizability333ability to correctly solve tasks on instances longer than those seen during training. Specifically, we show (Section 3) that theory predicts transformers to have a uniqueness bias (i.e., unique-token induction is easier than its non-unique counterpart) (Huang et al., 2025; Zhou et al., 2024); but for it to not have a directional bias (i.e., that forward and reverse variants are equally hard). Theory also predicts that within non-unique induction, retrieving is easier from the first occurrence of the query character than from the last one.
However, given the impressive capabilities of transformer-based LLMs, it is reasonable to posit that the massive scale of both such models and of pretraining might overcome these limitations. For instance, high prevalence of non-unique copying in natural language might lead transformers to learn specialized circuits or components for handling such situations. Here, we directly test whether these architectural limitations remain relevant to pretrained LLMs. We study 8 pretrained LLMs from Llama-3 and Qwen2.5 families (Grattafiori et al., 2024; Yang et al., 2024). Our findings show that large-scale pretraining produces a directional bias (Section 4.1). In addition to showing our results on synthetic settings, we show how these limits might surface and the reliability risks they pose for practitioners in more natural settings as well (Section 4.2). While targeted fine-tuning can even out the directional bias born out of pretraining, we find that the fundamental architectural uniqueness bias still binds pretrained transformers (Section 4.3).
2 Background, Notation & Definitions
Task Description.
We examine two primary task types, Retrieval and Copying, each instantiated under Unique and Non-unique conditions, as illustrated in Figure 1. In Retrieval, the sequence has the form $X=\langle bos\rangle x_{1},\dots x_{N-1}\langle sep\rangle x_{N}$, where each $x_{i}\in$ some alphabet $\Sigma$ and $x_{N}=q$ is the query token. The query $q$ appears $t$ times in the context, indexed as $q_{1},\dots q_{t}$ with $1\leq q_{1}<\dots<q_{t}\leq N$. Transformer has to predict a single token as the output given the input. For the Unique condition, we define UL (Unique Induction Left) and UR (Unique Induction Right) respectively when the output expected is the token preceding or following the query’s single occurrence, i.e., $x_{q_{1}-1}$ and $x_{q_{1}+1}$. Under the Non-unique condition, NLFirst (Non-unique Left First) and NRFirst (Non-unique Right First) use the first occurrence to predict $x_{q_{1}-1}$ and $x_{q_{1}+1}$ as the output, while NLLast (Non-unique Left Last) and NRLast [^4] The NRLast variant is conceptually equivalent to the Flip-Flop task of Liu et al. (2024a). (Non-unique Right Last) use the last occurrence to produce $x_{q_{t}-1}$ and $x_{q_{t}+1}$ respectively.
For Copying, the sequence is $X=\langle bos\rangle x_{1}\dots x_{N}\langle sep\rangle$ and the expected output is a continuation $x_{N+1}\dots x_{2N}$. The decoder is expected to replicate the latter segment based on the first. For Unique conditions all of the tokens in $X$ are unique. UF (Unique Forward copying) and UB (Unique Backwards copying) denote forward and reverse copying of the input segment. For Non-unique settings there is no uniqueness constraint on the $x_{i}$ values and NF and NB also denote the forward and reverse copying of the input segment.
Figure 1: Overview of our task variants (formal definitions in Section 2). Retrieval: A ‘#’ marks the separator; the token that follows is the query that may appear once (unique) or multiple times (non-unique). With non-unique queries we return the token immediately left/right of either the first or the last occurrence, creating 6 sub-tasks in total. Copying: We want to model to copy the context which consists of either only unique tokens or repeated tokens in the forward or reverse direction. Tasks in green lie in C-Rasp[pos] and length-generalize; those in red do not (proofs in Section 3).
Induction and Anti-Induction Heads.
Induction heads were characterized by Elhage et al. (2021) as part of a two-head circuit present in autoregressive transformers
Figure 2: Illustration of the induction and anti-induction Circuits.
and were later shown to be critical to performing in-context learning (Olsson et al., 2022).A previous-token head in layer $\ell$ writes the embedding of the immediately preceding token into the residual stream. An induction head in some layer $\ell^{\prime}\!>\!\ell$ then queries for positions whose stored “previous token” equals the current token and copies the corresponding value vector forward, enabling the model to predict the token after a repeated bigram. Our UR retrieval as well as the UF copying task setups are not novel and are meant to test this induction circuit ability and compare with our other task setups.
We hypothesize that, by symmetry, an anti-induction circuit should in principle exist and be able to copy information backwards. Here, the same previous-token head caches the token immediately to the left, but is paired with an anti-induction head that looks back to the earlier occurrence of the same token. Because that earlier token received the value of its own predecessor from the previous-token head, the anti-induction head can retrieve that predecessor, allowing the model to predict one step to the left. Such a two-head motif should thus enable exact solutions to our UL and UB tasks. Retrieval heads of Wu et al. (2025) subsume both induction and anti-induction heads (see Appendix E for discussion).
3 Theoretical Length Generalization Guarantees for Our Tasks
We derive theoretical predictions for transformers’ length-generalization abilities on the retrieval and copying tasks introduced above, using the framework of Huang et al. (2025), which links length-generalizability to expressiveness in C-Rasp (Yang and Chiang, 2024). Of particular interest is C-Rasp (Yang and Chiang, 2024), which restricts the usage of positional information and arithmetic operations. Huang et al. (2025) further added positionally-aware primitives to C-Rasp to construct the language C-Rasp[pos] [^5] Referred to as C-Rasp[local, periodic] in Huang et al. (2025). See Appendix A.1.. Doing so, they established–both theoretically and empirically–a tight link between the existence of a C-Rasp[pos] program for a task and the ability of decoder-only transformers (with absolute or no positional encodings) to length-generalize on that same task: in particular, if C-Rasp[pos] program exists, a decoder-only model provably generalizes to longer inputs under a specific formal model of training (Theorem 7 in Huang et al. (2025)). More precisely, this guarantee applies to APE Transformers under an idealized training procedure, where all training data up to some maximum input length $N$ are available and the exact minimizer of a regularized loss is found. Under these specific conditions, the existence of a C-Rasp[pos] program for that task guarantees length generalization to larger lengths. While theory has only been worked out for this specific setup, empirically we observe converging results across different PE types and realistic SGD-based training (as shown in Huang et al. (2025) as well as in the experiments in Appendix B). As a converse to this length generalization guarantee, extensive experimental evidence in Huang et al. (2025) suggests that, if no C-Rasp program exists, such models fail to generalize to longer inputs. [^6] While Huang et al. (2025) focuses on absolute (APE) or no positional encodings (NoPE), our empirical results show that the conclusions also match the behavior of Rotary Positional encodings (RoPE) (See Appendix B). Full theoretical understanding for RoPE length generalization remains an interesting problem. One interesting aspect is that, while training RoPE for UF and UB did not show perfect generalization, we did obtain perfect generalization when instead autoregressively decoding from models trained for UR/UL. In APE, we obtained perfect generalization when training directly for UF and UB, in line with C-Rasp[pos] expressivity.
We therefore adopt the following strategy for providing length generalizing guarantees across our retrieval and copying tasks:
[…]
We start by considering the “Right” versions of our Retrieval tasks. In the unique case (UR), a transformer can solve retrieval using a standard induction circuit (Figure 2): an induction head retrieves the symbol that followed the unique occurrence of the query. These operations can be simulated in C-Rasp[pos] (Lemma 4), ensuring a theoretical length generalization guarantee. In the non-unique case, expressibility – and hence the length generalization guarantee – varies:
Theorem 1.
NRFirst is expressible in C-Rasp[pos]. NRLast is not expressible in C-Rasp[pos].
The proof is in the Appendix (Lemmas 5 and 6). An extension of the induction circuit performs NRFirst: first, at each $x_{i}$, an attention head checks whether this symbol has appeared previously. Given a query $w$, an induction circuit then retrieves the symbol immediately following the (unique, if any exist) first appearance of $w$. A C-Rasp[pos] program formalizes this (Lemma 5). On the other hand, the inexpressibility of NRLast is proven by reduction to the regular language $(a|b|e)^{*}ae^{*}$, provably not in C-Rasp (Lemma 6). Regarding copying tasks, Huang et al. (2025) shows that UF is in C-Rasp[pos], while NF is not. The construction for UF again uses an induction circuit.
So far, we have considered “right” and “forward” versions—in fact, in all tasks, C-Rasp[pos] expressibility is equivalent for the left (backward) settings:
Theorem 2.
Across all tasks, there is no C-Rasp[pos] expressibility difference between R vs L versions, and F vs B versions.
That is, C-Rasp[pos] expressivity is the same for UL and UR, the same for UF and UB, etc. (Corollary 9). All tasks are annotated in Figure 1 with their C-Rasp[pos] expressibility. Experiments in Appendix B show that transformers trained from scratch on these tasks exhibit length generalization as predicted by the theoretical results.
Figure 3: In-context accuracy for Llama‑3 70B and Qwen2.5‑32B across all our tasks averaged over three seeds. Across all settings, lengths, model size, and task type, we observe a Directional Bias: Retrieving the token to the left of the query token is always more difficult compared to the one to the right, provided all other things are constant. Similarly, copying in the forward direction is easier than copying backwards. Detailed prompts, similar performance graphs on other models (including instruction-tuned variants) are in Appendix C.
4 Experiments
4.1 Eliciting Abilities via In-Context Learning
We evaluate pretrained completion as well as instruction-tuned variants of the following models: Llama3.1--8B, Llama3.1--70B, Qwen2.5--7B, and Qwen2.5--32B. For retrieval as well as copying, each model is evaluated across various prompt variations. The prompt variations relate to the few-shot example size (examples are equal in length to the test string or are smaller but clearer) and the instruction templates (details on prompt templates in the Appendix C.1) with the templates adjusted for completion as well as instruction-tuned models. All experiments use 1,500 test strings for each length 10,20,30,40,50,100{10,20,30,40,50,100} generated under three independent seeds. We cap the input string length at $100$ tokens to maintain a bounded vocabulary comparable to the unique setups; see Appendix C.4 for results at other lengths. The test strings contain spaced letters (including diacritics) and digits, to ensure that every character gets tokenized as an independent token and thus, the required uniqueness / repetitions can be maintained within the context, whenever required. Few‑shot prompts contain $k=5$ demonstrations, reserving one slot whenever an explicit explanation is required by the template. For each task that belongs to the same group i.e. for Unique (UR, UL) and for Non-Unique (NRLast, NRFirst, NLLast, NLFirst), we have the same strings on which we test the in-context ability with different expected answers to avoid the confound of different datapoints.
Directional Bias: Pretraining induces a left-right asymmetry. We found a persistent directional bias in all of the pretrained models we tested (Figure 3). For each task, retrieval or copying, the models consistently performs better on rightward tasks (UR, NRFirst, NRLast, UF, NF), across models, prompt templates, and sequence lengths. The drop in performance for the leftward tasks cannot be attributed to lexical cues such as the words “right” or “forward” (explored in earlier work (He et al., 2025)) because the performance asymmetry persists even when we strip all natural‑language instructions and provide only input–output pairs for the model to infer the task in-context. C-Rasp[pos] predicts symmetric performance in transformers trained from scratch within a task family (Theorem 2). We further find that models trained from scratch do not exhibit such a left-right asymmetry, confirming it is not inherent to the architecture (Appendix B). We therefore hypothesize that pretraining instills this bias because induction is a more natural task than anti‑induction and thus more induction head circuits get formed during pretraining as opposed to anti-induction circuits; Sections 4.3 and 4.4 probe this explanation in more detail.
Uniqueness Bias: Unique tasks are easier to elicit than non‑unique tasks. Unique retrieval and copying setups both perform relatively well in rightward / forward direction. In contrast, in the non-unique settings, eliciting correct retrieval in any direction is challenging across task setups. We observe an interesting difference to what is expected for transformers trained from scratch: First, the underperformance of NRFirst in pretrained models does not align with what is expected for transformers trained from scratch (Theorem 1). Copying, by contrast, is easier to elicit than retrieval, even though copying requires predicting many tokens and hence offers more opportunities for error. Remarkably, Non‑unique‑forward (NF) remains near perfect throughout. As this experiment only considers input lengths up to 100 for comparable results vis-à-vis the unique settings, an interesting question is how copying performs at longer lengths. The next section therefore stress‑tests in‑context copying on longer paragraphs, seeking to reveal if the theoretically predicted difficulty of non-unique copying (NF) appears in realistic settings.
Takeaway: Pretraining introduces a Directional Bias favoring retrieving and copying in the forward/right direction over the backwards/left. Pretraining also selectively amplifies the abilities of a transformer, as evident from the near perfect performance on copying repeated tokens, and poor performance on certain retrieval tasks, even though the latter are better learnable in a length-generalizable way for the architecture.
4.2 Testing Uniqueness Bias and Directional Bias in Natural Settings
Testing Uniqueness Bias by Copying Lorem Ipsum Paragraphs.
In Section 4.1, forward copying showed near-perfect performance irrespective of uniqueness (UF and NF). We now ask whether the theoretically predicted asymmetry between unique and non-unique copying still impacts LLM behavior, by having models reproduce longer, naturalistic but semantically neutral text. We here do not focus on evaluating exact copying accuracy, instead focusing on understanding the nature of LLM mistakes.
Figure 4: Failures in accurate copying of Lorem Ipsum paragraphs are associated primarily with ambiguous transition indices.
While entirely incorrect outputs are easily detectable, subtle differences in nearly accurate copies represent the most challenging errors for users to identify. Therefore, we focus exclusively on the cases where the generated output is within 75% of the length of the expected output, targeting scenarios where the model clearly understands the copying task and produces plausible copies of the input. To confirm that our prompting setup is valid, we evaluate the models on copying exact spans from recent arXiv papers (April 2025), finding perfect performance.
We design our copying benchmark to suppress semantic cues that could help the model break ambiguous copy chains. Specifically, each input consists of a randomized Lorem‑Ipsum–style paragraph. Such content is unlikely to have appeared verbatim in the training corpus. Each paragraph has close to 500 tokens. See Appendix C.5.1 for results at longer lengths (upto 5k tokens).
We classify each token in the input as either unambiguous (uniquely predictive of the subsequent token) or ambiguous (potentially followed by multiple subsequent tokens and therefore is not perfectly predictive of the subsequent token). We perform token-level alignments using the Needleman-Wunsch algorithm (Needleman and Wunsch, 1970), to detect token-level insertions, deletions, and substitutions. What we are left with is an alignment sequence consisting of 4 kinds of entries: match, insert, delete and substitute. We group consecutive entries in this alignment list if they are of the same type and call them aligned spans if the entries consist of only ‘match’ and misaligned spans if the entries grouped consist of any of the other types. We refer to indices where an aligned span finishes and a misaligned span starts as transition indices. These transition indices mark critical points where an error in copying started propagating during autoregressive generation. It should be noted that, even if there are just a few such misaligned spans, they can each be very long, especially for the ones consisting exclusively of ‘insert’. Such hallucinated insertions can at times snowball into a long error chain from where the model never reorients itself to produce a correct continuation (Examples in Appendix C.5).
We categorize the transition indices as unambiguous or ambiguous based on the nature of the token at that index. Since an induction head would suffice to correctly predict the subsequent token following an unambiguous token, we hypothesized that transition indices would never be unambiguous. Conversely, ambiguous tokens were expected to be the source of where all of the misaligned error span began, as it has similarities to the failure modes expected in NF (Non-unique forward copying). Empirical results (Figure 4) confirm this hypothesis. Across all model variants tested (sizes, families, completion as well instruction tuned models), a consistent pattern emerges, the transition indices almost unilaterally always are ambiguous and cause the ‘glitches’ we see in the final copied output. While these glitches appear across models, their exact nature appears to be random. We did not find any pattern except for the fact that they were tied to ambiguous tokens.
Testing Directional Bias in Git Commit History: Direction‑Aware Copying.
While the Lorem Ipsum setup shows how glitches in forward copying might cause reliability issues, we now show how failures in backward copying impact real world tasks. We use a compact Git history manipulation task to exemplify when the right-over-left asymmetry seen in Section 4.1 surfaces in practice. Given a truncated git log (newest→oldest commit hashes), the model must list commits that need revertinggit revert(forward order) or must be cherry‑pickedgit cherry-pick(reverse order) onto a release branch.
Figure 5: Git Commit History Manipulation also has the forward vs backward asymmetry seen in Section 4.1.
Llama‑3 models reproduce the forward list almost perfectly but accuracy drops when the order flips (Figure 5) – mirroring the asymmetry seen in the synthetic experiments (Section 4.1). Qwen models are not able to do the task at all, hence we exclude them from this analysis. The reversal presented here does not happen here at the token level, instead at the commit order level and is thus in principle more similar to a generalized copying setup. However, the theoretical results and expectations transfer to the generalized scheme (See Appendix for a detailed discussion A.2.1).
Takeaway: Even though pretraining selectively amplifies certain abilities of the transformer, fundamental limitations inherent to the architecture have a marked impact even after the pretraining and need to be kept in mind while deploying these models in real-world tasks.
4.3 Fine-tuning Eliminates the Directional Bias, But Not the Uniqueness Bias
Our experiments have uncovered three failure modes: (i) a Directional Bias in pretrained models, (ii) poor performance on certain retrieval tasks that are provably length-generalizable, and (iii) the Uniqueness Bias: sporadic “glitches” on copying tasks that are not length-generalizable. Each of these could stem either from architectural limitations or from inductive biases imparted during pretraining. To disentangle the two, we carry out task-specific supervised fine-tuning and see which failures are removed. Since we wish to know whether the model has truly learned a length-generalizable solution, we evaluate on test sequences twice as long as those seen in training, following Huang et al. (2025). A fully correct model should therefore achieve $100\%$ accuracy on this test set; even a single error then flags a reasoning flaw that could undermine reliability in real-world applications, an evaluation scheme similar to that of Liu et al. (2024a).
Figure 6: On the out-of-distribution dataset, tasks belonging to C-Rasp[Pos] (UL, UR, UF, UB, NLFirst, NRFirst) retain perfect accuracy, whereas NLLast, NRLast, NF, and NB, which are provably not in C-Rasp[Pos], suffer a drop in performance. The Directional Bias within task setups is largely absent from the fine-tuned models.
Experimental Setup. We fine-tune a pretrained language model —GPT–2 1.5B (APE) [^7] https://huggingface.co/karpathy/gpt2_1558M_final4_hf— on each retrieval task plus unique forward and backward copying. Fine-tuning uses the standard next-token objective but masks the loss to the answer span only, mimicking inference, i.e., for retrieval tasks, just the answer token, and, for the copying tasks, all tokens after the EOS delimiter in the input, indicating the end of the input sequence. To ensure that APE models can generalize on longer positions than those seen during training, we apply the positional offset trick of Huang et al. (2025): for every training input a random constant $\Delta\in[0,200-\ell_{min}]$ is added to all position indices, ensuring that all the positional embeddings are trained. All models see training examples whose input lengths are sampled uniformly from $\ell_{\min},100]$, where $\ell_{\min}$ is the length of the shortest well-formed instance of the task. We confirmed that, prior to this fine-tuning, the base model’s performance on each task was effectively zero. Details of hyperparameters and datasets are in Appendix D.
Our testbed setup for all of our tasks closely follows the setup of Bhattamishra et al. (2020); Huang et al. (2025), with an in-distribution bin (i.e., lengths $\ell_{min},100]$) and an out-of-distribution evaluation bin (i.e., lengths $[101,200]$). Similar to their setup, we report accuracy on our test sets only on seeds where training loss converges to 0. For our out-of-distribution evaluation bin on Retrieval tasks, we enforce that all instances of the query token $q$ appear exclusively in the first half of the context $\mathcal{C}$. This way distance between the relevant tokens to which the model needs to attend increases, thereby rigorously testing the model’s ability to generalize to longer contexts. In such settings, superficial heuristics should be insufficient, and thus we expect that only models with robust length generalization can perform reliably. We find that C-Rasp[pos] perfectly tracks length generalization results in our fine-tuning experiments: Every task which belongs to C-Rasp[pos], and is thus predicted to length-generalize, does so, whereas tasks not in C-Rasp[pos] do not length-generalize. Additionally, the Directional Bias of pretrained models disappears with fine-tuning, confirming the fact that it was an artifact of pretraining. We next seek to understand what leads to the development of this asymmetry by analyzing induction and anti-induction circuits (defined in Section 4.4) both in pretrained models, where this asymmetry exists, and in our fine-tuned models, where it doesn’t.
4.4 Source of the Directional Bias: A Mechanistic Perspective
We hypothesize that the source of directional bias in pretrained models lies in the difference between the presence and strength of induction and anti-induction circuits as described in Section 2. We posit that, in pretrained models, anti-induction heads are less common, but that fine-tuning can boost these, removing the asymmetry. We use Unique copying (UF and UB) setups to evaluate our hypothesis.
(a) Fine-tuning strengthens the relevant heads
(b) Induction and Anti-Induction heads are causal
Figure 7: (a) Top-10-percent attention heads ranked by their attention score, i.e. the mean attention weight that target-string tokens allocate to either the same source token (anti-induction) or the next source token (induction). Dashed lines: pretrained checkpoints; solid lines: fine-tuned. Shaded bands cover source lengths 3–10. Fine-tuning eliminates the pretraining imbalance; for forward copying (UF) fine-tuning strengthens induction heads, for backwards copying (UB) fine-tuning boosts anti-induction heads. (b) Patching experiment that ablates each family of heads while measuring task accuracy. For UF, removing induction heads (red $\rightarrow$ green) is fatal, whereas ablating anti-induction heads is inconsequential (red $\rightarrow$ blue); the roles reverse for UB. This holds across models & sequence lengths confirming the causal status of the two circuits.
Induction and anti-induction heads are causal in both fine-tuned and pretrained models. We conduct a patching experiment to confirm that removal of induction heads hurts performance in the UF task, while removal of anti-induction heads hurts performance in the UB task. Indeed, both in fine-tuned and pretrained models, removing induction heads is catastrophic for UF and has a negligible effect in UB, and removing anti-induction heads cripples backward copying while leaving forward copying intact. On fine-tuned checkpoints, the effect is amplified to the extreme: removing relevant heads drops the accuracy to zero, while removing another type of heads has no effect (Figure 77(b)). Thus, induction and anti-induction heads have a causal effect on model performance (Details of patching experiment in Appendix E). Additionally, the induction and anti-induction heads get amplified via fine-tuning (Figure 77(a)). Fine-tuning strengthens the heads that matter: induction heads gain attention weight for forward copy, anti-induction heads for backward copy.
Takeaway: The directional bias in pretrained transformers can be explained by the difference in strength and prevalence of induction vs. anti-induction heads. Fine-tuning can be used to amplify either of these heads and thus remove the asymmetry.
5 Discussion
Related Work.
Length generalization has been studied extensively (Anil et al., 2022; Abbe et al., 2023; Zhou et al., 2024), including attempts to improve it using scratchpads (Abbe et al., 2023; Hou et al., 2024) or specific positional embedding schemes (Press et al., 2022; He et al., 2024; Cho et al., 2024). Transformer limitations: besides copying and retrieval, transformers struggle with problems including sensitive functions (Hahn and Rofin, 2024) and state tracking (Merrill and Sabharwal, 2023). These limitations can be related to failures in LLMs on tasks such as multiplication (Frieder et al., 2024; Satpute et al., 2024; Amiri et al., 2025) and maintaining state information (Toshniwal et al., 2022; Kim and Schuster, 2023; Merrill et al., 2024; Zhang et al., 2025). Induction heads have been studied extensively since Elhage et al. (2021) and their existence has been linked to in-context learning and copying (Olsson et al., 2022; Chen et al., 2024a; Singh et al., 2024; Crosbie and Shutova, 2025). Retrieval heads (Wu et al., 2025) generalize this notion by conditionally copying information from arbitrary positions, subsuming our definitions of induction and anti-induction heads.
Takeaways.
While failures in length generalization in copying and retrieval tasks may not affect typical user interaction, they pose reliability concerns in specialized applications. For instance, AI coding assistants like GitHub Copilot (Chen et al., 2021) routinely handle non-semantic identifiers (commit hashes, variable names, function names). Glitches in copying such strings could lead to bugs in the deployment of critical code and would therefore be highly undesirable. Pretraining of LLMs has been shown to impart biases, be it for specific tokens or specific entities (He et al., 2025) and in-context performance can be brittle based on the choice of tokens used. This can lead to reliability risks in domains requiring high precision such as medicine (Li and Chong, 2024; Bélisle-Pipon, 2024). In contrast, the left-right asymmetry we uncover is indicative of a latent ability that is difficult to elicit in pretrained models without explicit fine-tuning. We ask practitioners to be aware of the existence of such biases and consider task-specific fine-tuning to eliminate this kind of asymmetry if it is considered undesirable for a downstream application.
On the other hand, glitches in copying large semantically neutral text (SHA keys, DNA sequences) would fall under the umbrella of repeated copying. To achieve complete reliability and compensate for persistent architectural limitations, practitioners might need to resort to external tool usage in LLMs. Similar to how tool usage is used to compensate for mathematical weaknesses in LLMs (Gao et al., 2023; Chen et al., 2024b; Gou et al., 2024; Wang et al., 2024), designing principled policies that decide when to use tools like programmatic copy-pasting could be an interesting research direction.
Limitations.
We focus on two open LLM model families, not covering the entire model landscape, especially closed source systems or models based on recurrent architectures, which might show different sets of abilities or limitations (Sarrof et al., 2024; Grazzi et al., 2025; Siems et al., 2025). Second, due to resource limitations, fine-tuning experiments in Section 4.3 are limited to a single, 1.5B sized model. We also do not analyze the underlying pretraining data of the models to get an estimate of what the correct in-distribution bin should be for any of our models. Larger contexts should in principle push the out-of-distribution boundary to higher lengths as the in-distribution size increases. It remains an open question whether length generalization would even matter for larger models with a huge in-distribution context size. Finally, our theoretical results linking C-Rasp[pos] to length-generalizability in Huang et al. (2025) have only been proven formally for APE, and applicability to RoPE has only been tested empirically.
6 Conclusion
We asked whether large‐scale pretraining can erase the transformer’s native length‐generalization limits, using a foundational family of retrieval and copying tasks as testbed. We find that pretraining boosts certain abilities, but transformers remain bound by the same length-generalization limits they were “born” with: Pretraining amplifies induction circuits, making right-/forward-oriented retrieval and copying notably easier, but does not overcome the difficulty inherent to non-unique retrieval and copying. Fine-tuning can rebalance induction and anti-induction heads and restore full, theory-aligned generalization, but only by explicitly teaching the model its missing capacities. Our results thus suggest that, while large-scale pretraining brings remarkable capabilities, it cannot fundamentally rewrite the architecture’s core inductive biases.
Acknowledgements
We thank anonymous reviewers for their encouraging and constructive feedback. This research is funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102 “Information Density and Linguistic Encoding”; Project-ID 471607914 – GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action”; and Project-ID 389792660 – TRR 248 “Foundations of Perspicuous Software Systems”. We would like to thank Mark Rofin, Blerta Veseli, Brian DuSell, Ekaterina Shutova, Kayo Yin, Xinting Huang, Yifan Wang for discussions and feedback on the draft.
Contributions
MJ contributed to Sections 4.1, 4.2, 4.3, scaling experiments, paper refinement and the Appendix. YV contributed to Sections 4.1, 4.2, 4.3, 4.4, and paper refining. YS contributed to Sections 4.1, 4.2, the Appendix, scaling experiments, paper drafting and refining. SB contributed to Section 4.4. VD and EP provided feedback, supervision, and refined the paper. MH supervised the project and contributed to Section 3, paper drafting, and refining.
Appendix A Detailed Theoretical Proofs
A.1 C-Rasp[pos] as a Framework for Predicting Length Generalization
C-Rasp[pos] (pos for “positional information”) was introduced as C-Rasp[periodic,local] by Huang et al. [2025] [^8] We use a shorter name just for convenience. as a formalization of Rasp-L and the Rasp-L conjecture of Zhou et al. [2024]. It extends C-Rasp [Yang and Chiang, 2024] by incorporating positional information. Huang et al. [2025] connected expressibility in C-Rasp[pos] to length generalization for transformers both theoretically and empirically:
  1. 1.
    They proved that transformers, when trained with an idealized model of learning, generalize on problems expressible in C-Rasp[pos].
  2. 2.
    Across a battery of algorithmic and formal language problems, they showed that problems expressible in C-Rasp[pos] show length generalization when transformers are trained with SGD.
  3. 3.
    For problems in this battery that are not expressible in C-Rasp[pos], they showed that length generalization was empirically unsuccessful.
These results unify and explain a variety of prior results; for instance, the link between C-Rasp[pos] and length generalization explains why transformers show persistent glitches in FlipFlop (empirically observed by Liu et al. [2024a]), it explains why length generalization on copying is difficult in the presence of repetition [Zhou et al., 2024, Jelassi et al., 2023], and why length generalization on addition is difficult [Zhou et al., 2024]. Importantly, C-Rasp[pos] offers a single framework for understanding length generalization, with well-understood methods for understanding expressivity. It is more fine-grained than various other methods for understanding transformers’ expressivity: For instance, all tasks considered in Zhou et al. [2024], Huang et al. [2025] are in $TC^{0}$ and in principle expressible by transformers, but length generalization varies substantially in line with C-Rasp[pos] expressivity. Thus, C-Rasp[pos] provides a more fine-grained perspective by focusing on length generalizability.
C-Rasp[pos] can thus be viewed as a formal model of the space of problems that transformers can represent across different input lengths and are likely to length-generalize on. transformers can in principle represent problems outside of C-Rasp[pos], but these are then likely to not show length generalization because representations at increasing lengths will require increases in model size or parameter norm – if learning tends to find simpler solutions, one will expect length generalization to fail [Zhou et al., 2024, Huang et al., 2025]. For instance, one can construct a transformer performing copying over strings of any fixed length $N$, e.g. by hard-coding matching pairs of positions [Bhattamishra et al., 2024], but any construction that works across all lengths up to $\leq N$ will require a growth of model size (in a sense made formal by Huang et al. [2025]) as $N$ increases, i.e., length generalization is unlikely.
As Huang et al. [2025] formalizes the Rasp-L conjecture of Zhou et al. [2024], we expect that the same conclusions would likely be reached using the Rasp-L language from Zhou et al. [2024]; the advantage of C-Rasp[pos] as compared to Rasp-L is that its expressiveness is provably understood (including impossibility proofs), and is theoretically linked to length generalization (Theorem 7 in Huang et al. [2025]).
For self-containedness, we recapitulate the operations used in C-Rasp[pos] here:
Definition 3.
Let $\Sigma$ be an alphabet.
Let $\Phi$ be a set of unary relations $\phi:\mathbb{N}\rightarrow\{\bot,\top\}$ where each $\phi$ satisfies
  • (Periodicity) for some $\Delta>0$, it holds that $\phi(i)=\phi(i+\Delta)$ for all $i$
Let $\Psi$ be a set of binary relations $\psi:\mathbb{N}\times\mathbb{N}\rightarrow\{\bot,\top\}$ here each $\psi(i,j)$ satisfies [^9] These properties were referred to as translation invariance and locality in Huang et al. [2025].
  • (Depends only on Distance) $\psi(i,j)$ only depends on iji-j
  • (Restricted to Bounded Distance) $\{i:\psi(i,j)=\top\}$ is finite, for any $j\in\mathbb{N}$
Here, $\bot$ stands for “false” and $\top$ stands for “true”.
A C-Rasp[pos] program $P$ is defined as a sequence $P_{1},\ldots,P_{k}$ of operations, each of the following types:
Boolean-Valued Operations
Initial $P(i):=Q_{\sigma}(i)$
           for $\sigma\in\Sigma$
Boolean $P(i):=\lnot P_{1}(i)$
$P(i):=P_{1}(i)\land P_{2}(i)$
Constant $P(i):=\top$
Positional $P(i):=\phi(i)$
           for $\phi\in\Phi$
Comparison $P(i):=C_{1}(i)\leq C_{2}(i)$
Count-Valued Operations
Counting $C(i):=\textsc{{\#}}\left[j\leq i,\psi(i,j)\right]\;P(j)$
           for $\psi\in\Psi\cup\{\top\}$
Conditional $C(i):=P(i)\;\mathbf{?}\;C_{1}(i)\;\textbf{:}\;C_{2}(i)$
Addition $C(i):=C_{1}(i)+C_{2}(i)$
Subtraction $C(i):=C_{1}(i)-C_{2}(i)$
Constant $C(i):=1$
The most important construct is the Counting operation, which returns an integer indicating for how many positions $j\leq i$ both $P(j)$ and $\psi(i,j)$ hold. In the special case where $\psi=\top$ (i.e., a predicate that always returns “true”), the operation just counts how often $P(j)$ holds for $j\leq i$. Another important operation is the Initial operation, where $Q_{\sigma}(i)$ is true if and only if the $i$-th symbol in the string equals $\sigma$. We refer to Huang et al. [2025] and Yang and Chiang [2024] for detailed definitions of the semantics.
The most important example for a function $\psi\in\Psi$ is the “predecessor” predicate checking if j=i1j=i-1:
$\psi(i,j)=\begin{cases}\top&\text{if }j=i-1\\ \bot&\text{else}\end{cases}$ (1)
We will follow Huang et al. [2025], Yang and Chiang [2024] in using, as syntactic sugar, $\exists[j\leq i,\psi(i,j)]P(j)$ as a shorthand for $1\leq\#[j\leq i,\psi(i,j)]P(j)$, and $\textsc{{\#}}\left[j\leq i\right]\;P(j)$ for $\textsc{{\#}}\left[j\leq i,\top\right]\;P(j)$.
In this paper, we are interested in tasks that, given a prefix, produce a token: either the overall output symbol in a retrieval task, or the next output symbol when copying a string. To formalize this, we assume that every program contains Boolean-valued operations named $\text{NEXT}_{a}(i)$ for each $a\in\Sigma$ such that $\text{NEXT}_{a}(i)$ holds if and only if the correct next symbol is $a$. Thus, we say that a task is expressible in C-Rasp[pos] if there is a program defining $\text{NEXT}_{a}(i)$ for each $a\in\Sigma$ such that, for each $i$ in the span of the output (a single position for a retrieval task, the full output string for a copying task), $\text{NEXT}_{a}(i)$ holds (i.e, returns “true”) at the final position $i$ if and only if the correct next symbol is $a$. Generation stops when none of these operations return “true”.
A.2 Expressiveness Results
We show:
Lemma 4.
UL, UR are expressible in C-Rasp[pos].
Proof.
A UR program is given in Huang et al. [2025]; the UL program is analogous; overall:
C-Rasp[pos] program for UR, UL $\text{PRED}_{a}(i)$ $:=\exists[j\leq i,j=i-1]{Q_{a}(j)}$ (1) for each $a\in\Sigma$ $CBIGRAM_{ab}$ $:=\exists[j\leq i]{Q_{b}(j)\land\text{PRED}_{a}(j)}$ (2) for each $a,b\in\Sigma$ $\text{NEXT}_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land\text{CBIGRAM}_{\sigma a}(i)\right]$ (UR version) (3) for each $a\in\Sigma$ $\text{NEXT}_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land\text{CBIGRAM}_{a\sigma}(i)\right]$ (UL version) (4) for each $a\in\Sigma$
This program is a direct formalization of the (anti-)induction head circuit, and basic to the other C-Rasp[pos] programs constructed below. The first line checks for each $a\in\Sigma$, at each position $i$, whether position i1i-1 holds the symbol $a$. The second line checks, for each bigram $ab\in\Sigma\times\Sigma$ whether it appears at some position in the context up to the $i$-th position. The third line then states that, in UR, the correct next token is $a$ if and only if the bigram $\sigma a$ has appeared in the context, where $\sigma$ is the symbol at position $i$. The fourth line is the analogous instruction for UL: Here, $a$ is predicted if and only if the bigram $a\sigma$ appears in the context. ∎
Lemma 5.
NRFirst and NLFirst are expressible in C-Rasp[pos].
Proof.
We first show the construction for NRFirst; it builds on the construction from Lemma 4:
C-Rasp[pos] program for NRFirst $\text{ISLEFTMOST}(i)$ $:=\bigvee_{a\in\Sigma}\left[Q_{a}(i)\wedge\left(\#[j\leq i]{Q_{a}(j)}\right)\leq 1\right]$ (1) $\text{PRED}_{a}(i)$ $:=\exists[j\leq i,j=i-1]\left[Q_{a}(j)\wedge\text{ISLEFTMOST}(j)\right]$ (2) for each $a\in\Sigma$ $\text{CBIGRAM}_{ab}$ $:=\exists[j\leq i]\left[{Q_{b}(j)\land\text{PRED}_{a}(j)}\right]$ (3) for each $a,b\in\Sigma$ $\text{NEXT}_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land\text{CBIGRAM}_{\sigma a}(i)\right]$ (4) for each $a\in\Sigma$
where $\text{NEXT}_{a}(i)$ holds at the final position if and only if the desired completion is the symbol $a$. The first line in the program checks, at each position, if it is the leftmost representative of the symbol that it holds: That is, whether there is a symbol $a\in\Sigma$ that occurs at position $i$ but at no earlier position. For each $a\in\Sigma$, the second line checks, at position $i$, whether the preceding position i1i-1 holds the symbol $a$ – simulating the operation of a Previous-Token head. Additionally it also uses the computations made in the first line to check whether position i1i-1 was the leftmost representative with such a property. For each possible bigram $ab\in\Sigma\times\Sigma$, the third line checks whether it has appeared or not. Finally, the fourth line says that $a$ is the correct next token if and only if the current position holds a symbol $\sigma$ such that $\sigma a$ appears as a bigram in the context. This last line is the one that simulates the operation of the Induction head itself.
The program for NLFirst is very similar, it only differs in the fourth line:
C-Rasp[pos] program for NLFirst $\text{ISLEFTMOST}(i)$ $:=\bigvee_{a\in\Sigma}\left[Q_{a}(i)\wedge\neg\exists[j\leq i]{Q_{a}(j)}\right]$ (1) $\text{PRED}_{a}(i)$ $:=\exists[j\leq i,j=i-1]\left[Q_{a}(j)\wedge\text{ISLEFTMOST}_{a}(j)\right]$ (2) for each $a\in\Sigma$ $\text{CBIGRAM}_{ab}$ $:=\exists[j\leq i]\left[{Q_{b}(j)\land\text{PRED}_{a}(j)}\right]$ (3) for each $a,b\in\Sigma$ $\text{NEXT}_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{a}(i)\land\text{CBIGRAM}_{a\sigma}(i)\right]$ (4) for each $a\in\Sigma$
where $\text{NEXT}_{a}(i)$ holds at the final position if and only if the desired completion is the symbol $a$.
Lemma 6.
NRLast and NLLast are not expressible in C-Rasp[pos].
Proof.
We show this by reducing them to Lemma 36 in Huang et al. [2025]. Assume, for the sake of contradiction, that NRLast was expressible in C-Rasp[pos]. This program would discriminate between the two languages
$(w0|w1|i)^{*}w1i^{*}\langle sep\rangle w\text{\ \ \ \ \ \ \ \ \ \ \ \ (retrieving a ``1'')}$ (5)
and
$(w0|w1|i)^{*}w0i^{*}\langle sep\rangle w\text{\ \ \ \ \ \ \ \ \ \ \ \ (retrieving a ``0'')}$ (6)
Such a program could be turned into a program discriminating between the two languages
$(w0|w1|i)^{*}w1i^{*}$ (7)
and
$(w0|w1|i)^{*}w0i^{*}$ (8)
By a simple transformation of the alphabet, this in turn would amount to discriminating between the two languages
$(a|b|i)^{*}ai^{*}$ (9)
and
$(a|b|i)^{*}bi^{*}$ (10)
As described in the proof of Lemma 36 in Huang et al. [2025], a C-Rasp[pos] program discriminating between these two languages could be used to create a C-Rasp[pos] program for recognizing the language
$(a|b|i)^{*}bi^{*}b(a|b|i)^{*}$ (11)
which in turn is impossible by Lemma 35 in Huang et al. [2025]. The same reasoning holds for NLLast. ∎
Lemma 7.
UF, UB are expressible in C-Rasp[pos].
Proof.
This task can be viewed as iterated application of UL and UR, respectively, and can thus be done with the same program. The only complication is that (i) at the beginning of copying, at $\langle sep\rangle$, the program needs to retrieve the symbol that had followed $\langle bos\rangle$, and that (ii) generation should stop once $\langle sep\rangle$ has been generated a second time. These extra conditions can be encoded into C-Rasp[pos]. ∎
Lemma 8.
NF, NB are not expressible in C-Rasp[pos].
Proof.
Huang et al. [2025] (Theorem 12) show that NF is not in C-Rasp[pos] by showing that all problems expressible C-Rasp[pos] have logarithmic communication complexity in the model where Alice and Bob have access to the first and second halves of the string, respectively. Because NF has no sublinear communication protocol in this model, it is not in C-Rasp[pos]. The same argument applies to NB. ∎
We conclude from all lemmas in this section that there is no theoretical difference between left and right versions of our retrieval and copying tasks:
Corollary 9 (Repeated from Theorem 2).
Across all tasks, there is no C-Rasp[pos] expressibility difference between R vs L versions, and F vs B versions.
Proof.
Immediate from the lemmas in this section. ∎
A.2.1 Generalized reverse copying
Let the token alphabet be $\Sigma=\{\sigma_{1},\dots,\sigma_{z}\}\cup\{\mathord{\square}\}$ where $\mathord{\square}$ represents separators. $\Sigma_{\mathrm{w}}=\Sigma\setminus\{\mathord{\square}\}$ are all tokens that can be part of a word. Define the language $L=(\Sigma_{\mathrm{w}}^{+}\mathord{\square})^{*}\Sigma_{\mathrm{w}}^{+}.$ Every $x\in L$ can be written uniquely as $x=w_{1}\mathord{\square}w_{2}\mathord{\square}\cdots\mathord{\square}w_{n}$ with $w_{i}\in\Sigma_{\mathrm{w}}^{+}$ (i.e. each $w_{i}$ represents a word.) and also as $x=t_{1}\dots t_{m}$, where each $t_{i}\in\Sigma$ and is the token level representation of the string. We can think of copying this string $x\in L$ in 2 ways.
$\rho_{\mathrm{tok}}(t_{1}\ldots t_{m})=t_{m}\ldots t_{1},\qquad\rho_{\mathrm{word}}(w_{1}\mathord{\square}\cdots\mathord{\square}w_{n})=w_{n}\mathord{\square}\cdots\mathord{\square}w_{1}.$
$\rho_{\mathrm{tok}}(t_{1}\ldots t_{m})$ is the same as our backward copying setup, while $\rho_{\mathrm{word}}(w_{1}\mathord{\square}\cdots\mathord{\square}w_{n})$ more closely matches how a typical user would ever want to interact with an LLM to reverse a string if ever.
Corollary 10.
$\rho_{\mathrm{word}}$ is not expressible in $\mathsf{C\text{-}\textsc{Rasp}[pos]}$.
Proof sketch.
Theorem 5 states that Non‑Unique Copy backwards which is the same as $\rho_{\mathrm{tok}}$—lies outside $\mathsf{C\text{-}\textsc{Rasp}[pos]}$ by a communication‑complexity argument. To transfer this inexpressibility to $\rho_{\mathrm{word}}$ consider the following reduction.
Reduction.  Embed any token string $t_{1}\ldots t_{m}\in\Sigma^{*}$ into $L$ by inserting separators:
$t_{1}\mathord{\square}t_{2}\mathord{\square}\cdots\mathord{\square}t_{m}\;\in L.$
Because each “word” now has length $1$, applying $\rho_{\mathrm{word}}$ yields
$\rho_{\mathrm{word}}\bigl(t_{1}\mathord{\square}\cdots\mathord{\square}t_{m}\bigr)\;=\;t_{m}\mathord{\square}\cdots\mathord{\square}t_{1},$
and deleting the $\mathord{\square}$ symbols recovers $\rho_{\mathrm{tok}}(t_{1}\ldots t_{m})$. Hence an algorithm for $\rho_{\mathrm{word}}$ would give one for $\rho_{\mathrm{tok}}$ with constant overhead, contradicting Theorem 5. ∎
Intuition.
$\rho_{\mathrm{word}}$ reverses a list of chunks, whereas $\rho_{\mathrm{tok}}$ does the same when every chunk is a single token. When each word has length $1$ the two coincide, so solving word‑order reversal would solve Non-unique copy‑backwards as a special case—which creates a contradiction, and thus $\rho_{\mathrm{word}}$ should also not be solvable.
Appendix B From Scratch Training Experiments
Here, we validate theoretical predictions about the success of transformers in length generalization on all retrieval and copying tasks, by training small transformers from scratch. We particularly validate that results hold both for APE (as in GPT-2, as originally targeted by the theory of Huang et al. [2025]) and RoPE (as found in various other modern LLMs). We adopt the methodology from Huang et al. [2025] to demonstrate all our experiments when training transformers from scratch. Like Zhou et al. [2024], we sample independent training batches on the fly instead of using a finite-size training set. In contrast, each test set contains $2000$ samples that are sampled at the beginning of each experiment.
We train models with the default CausalLMLoss from the transformers library, the length of inputs is sampled uniformly from minimum up to maximum length in the specified range. This is true for training data where the range is $(l_{min},50)$ where $l_{min}$ is the minimum length possible for a given task (varies from 2 - 4, depending on the task variant and the minimum length of characters required to make the string valid). Our test sets have the size – $(l_{min},50),(51,100),(101,150)$. At each step, the model outputs the correct continuation – for retrieval the correct token to be retrieved, for copying the subsequent next token from the input sequence. The models are trained on a whole sequence of tokens. We train decoder-only transformer from scratch, using implementations from HuggingFace transformers for APE (GPT2LMHeadModel) and RoPE (LlamaForCausalLM). We train models for maximum 3030K steps with a batch size of $64$. We stop training early once the model’s accuracy reaches 100% on the in-distribution test set (the one in range $l_{min},50]$. The model is trained with a dropout rate of $0.0$, and we use AdamW, with a weight decay rate of $0.01$ (choices based on Huang et al. [2025]).
For experiments with APE, at training time, we add random offsets to position indices so that all position embeddings are trained (following Huang et al. [2025]). The offsets are sampled uniformly at random in the range $0,150-\ell_{curr}]$ where $\ell_{curr}$ is the length of the current string. For RoPE, such an offset is not required in training [^10] In fact, relative positional encodings such as RoPE are by definition invariant to the addition of a constant offset to the indices; instead a scaling factor of $32.0$ is set; the RoPE scaling type used was linear.
In preliminary experiments, we found that different model architectures, while achieving $100\%$ accuracy on in-distribution data, may perform differently on out-of-distribution data. To draw a conclusion about how the model performs on a problem in general, we determine the hyperparameters as follows: We consider configurations of {1, 2, 4} layers, {1, 2, 4} heads and model dimension of {16, 64, 256}, and learning rate of {0.001, 0.0001}. We sweep all the configurations by iterating over every combination and choose the one that achieves the highest accuracy on $[51,100]$ among those configurations whose accuracy on $l_{min},50]$ is 100%. When there are multiple such options, e.g., their accuracy on $[51,100]$ is 100%, the one with the simplest architecture is selected (when estimating complexity, we assume the following priority: number of layers >> number of heads >> model dimension). The final hyperparameters we used for each task are shown in Table 1 and 2. After we determine the hyperparameter configuration, we run the experiments with multiple random seeds and report the average accuracy of 3 successful runs (those runs where the model achieves 100% accuracy on in-distribution data).
Problem Model Size LR Max Steps
UL, UR, UF, UB 2 layer; 4 head; 64 dim 1e-3 30k
NLFirst, NRFirst 4 layer; 4 head; 64 dim 1e-3 30k
NLLast, NRLast, NF, NB 4 layer ; 4 head; 256 dim 1e-4 30k
Table 1: Hyperparameters for training with APE on transformers trained from scratch
Problem Model Size LR Max Steps
UL, UR, UF, UB 2 layer; 4 head; 64 dim 1e-3 30k
NLFirst, NRFirst 2 layer; 4 head; 64 dim 1e-3 30k
NLLast, NRLast, NF, NB 4 layer ; 4 head; 256 dim 1e-4 30k
Table 2: Hyperparameters for training with RoPE on transformers trained from scratch
UF, UB vs UR, UL:
While training UF and UB with RoPE, we did not observe perfect length generalization when directly training on the task, unlike what we observed with APE. However, we note that UF and UB can be thought of as repeated applications of UR and UL, for which we did obtain perfect length generalization even with RoPE – showing that length-generalizable RoPE transformers do exist for UF and UB. [^11] We leave to future research to determine whether this phenomenon reflects some deeper discrepancy between RoPE and APE, or superficial optimization difficulties. We thus use the RoPE transformers for UR and UL to model UF and UB. We take separate versions of transformers on UR and UL (without any separator before the query token) and autoregressively use the UR and UL models to generate a full continuation till an EOS is reached to get correct solutions for UF and UB. Therefore we format the input as: X = $\langle bos\rangle\langle sep\rangle x_{1}\dots x_{N}\langle eos\rangle\langle sep\rangle$. Upon seeing the separator, the model predicts $x_{1}$, and then $x_{2}$ and so on till we get to $x_{N}$, and then an $\langle eos\rangle$ is predicted. We do the same with UL to get predictions for UB.
Figure 8: Transformers trained from scratch with APE, perfectly aligns with C-Rasp[pos] predictions.
Figure 9: Transformers trained from scratch with RoPE, aligns with C-Rasp[pos] predictions, even though the length generalizability guarantees given by C-Rasp[pos] only apply to APE. It should be noted that the accuracy on UF, UB plotted here are not obtained by directly training, but by applying UR, UL repeatedly as described in the paragraph B. The bins here correspond to the evaluation bins of sizes $(l_{min},50),(51,100),(101,150)$).
Appendix C Additional experimental details and results
C.1 Prompting: In-Context Retrieval Templates (Section 4.1)
Table 3 below lists the three prompt factors and their possible values. A Cartesian product of all prompt settings across the 3 categories given below was taken to generate 20 prompt templates.
Table 3: Prompt Template Grid Factors
Factor Options
Separator sep, nosep
Few‑shot small, same
Template bare, simple_rule, simple_rule_expl., math_rule math_rule_expl.
Separator Examples
The input string is given in the following format based on the category.
  • SEP : $x_{1},x_{2}...x_{n-1}||x_{n}$
  • NOSEP : - $x_{1},x_{2}...x_{n-1}x_{n}$
Few‑Shot Pools
  • small — Examples drawn from a held‑out pool with lengths smaller than the input string $X$.
  • same — Examples sampled from the same evaluation set as the input string $X$ (excluding $X$ itself).
Instruction Templates
The different possible instruction templates for each of the tasks are given below. In every case, the prompt ends with the unresolved query line, prompting the model to emit the answer token.
bare
This template provides only the few-shot examples without any additional instructional or explanatory text. It requires the completion model to infer the task directly from the presented examples. Example for $UR$ :
simple_rule
The task is introduced through a concise English-language description defining the resolution rule. The idea of the setup is the test whether the model can follow explicit instructions based on linguistic clarity. The format across the tasks is as follows.
Each line is written as ‘context|query: target’.
The vertical bar ‘|’ separates the context from the query token.
All strings below follow this rule: {rule_simple}
{examples}
Table 4: Simple Rules for each of the tasks
Task Template
UL The answer is the token immediately to the left of the single instance of the query token
UR The answer is the token immediately to the right of the single instance of the query token.
NLLast When the query appears multiple times, the answer is the token just to the left of its last appearance.
NRLast When the query appears multiple times, the answer is the token just to the right of its last appearance.
NLFirst When the query appears multiple times, the answer is the token just to the left of its first appearance.
NRFirst When the query appears multiple times, the answer is the token just to the right of its first appearance.
simple_rule_explained
This template enhances the simple English rule format by providing an additional worked example explanation. It begins by stating the simple rule, then illustrating through an example how the rule is actually applied. Following the same, the model is presented with some few-shot examples to reinforce the rule understanding.
math_rule
This template explicitly formulates the task using formal mathematical definitions and notation. It introduces variables to clearly define the input string, context, query token, token indices, and continuation token. The mathematical notation precisely indicates conditions under which the continuation token should be selected, leaving no ambiguity regarding the required solution path. The template is as follows.
[…]
\$q\$ appears \$t\$ times at indices \$q_1, $\ldots$, q_t\$ (\$1 $\le$ q_1 $\le$ q_t $\le$ n-1\$).
[…]
Examples:
{examples}
Table 5: Math Rule Template
Task Template
UL $t=1$ and $x_{n+1}=x_{q_{1}-1}$
UR $t=1$ and $x_{n+1}=x_{q_{1}+1}$
NLLast $t>1$ and $x_{n+1}=x_{q_{t}-1}$
NRLast $t>1$ and $x_{n+1}=x_{q_{t}+1}$
NLFirst $t>1$ and $x_{n+1}=x_{q_{1}-1}$
NRFirst $t>1$ and $x_{n+1}=x_{q_{1}+1}$
math_rule_explained
Similar to the simple_rule_explained, this template begins with a detailed mathematical formulation of the task and is then followed by a worked-out example and then finally some few-shot set of strings.
C.2 Prompting: In-Context Copying Templates (Section 4.1)
Template variants.
For our In-context Copying task suite, we use three lightweight prompt styles. There is no Cartesian grid as in retrieval, as each of the prompts had statistically little variance, and eliciting this ability from our LLMs was not as challenging as in retrieval.
  • bare — few-shot examples only, no instructional text.
  • obey — a one‑line English rule (rule_simple) introducing the few‑shot block.
  • hint — a short hint (rule_hint) that highlights the input–output relation; mainly useful for backward copying.
Rule strings injected by each template.
The phrases referenced above are drawn from Table 6. Note that unique and non‑unique regimes share identical wording because the operation (copy vs. reverse) is independent of token repetition once the input is fixed.
Table 6: rule_simple phrases for the four copying tasks.
Task rule_simple
UF The output is exactly the same sequence as the input.
UB The output is the input sequence written in reverse order.
NF The output is exactly the same sequence as the input.
NB The output is the input sequence written in reverse order.
The corresponding rule_hint strings swap the lead‑in "The output is …" with "In every example the output …" to make the instruction less imperative.
Few‑shot sampler.
All copy prompts use the same held‑out k=5k{=}5‑shot pool (length L=5L{=}5, N=1500N{=}1500) that is disjoint from the evaluation set. Each prompt is built by picking five random lines from this pool, followed by the query built from the current test string.
End‑to‑end prompt bare template example (UF)
<bos> r m a j s : r m a j s <eos>
<bos> H w s x n : H w s x n <eos>
<bos> Q o F G J : Q o F G J <eos>
<bos> O a Y M F : O a Y M F <eos>
<bos> m I N r D : m I N r D <eos>
<bos> Z i b E B :
The unresolved line after is the model’s query; generation proceeds with greedy decoding ($T=0$).
C.3 Longer Vocabulary (Word-Level) Experiments (Section 4.1)
To ensure that the Uniqueness and Directional Biases observed in our primary character-level experiments are not simply artifacts of a synthetic setup, we conducted a parallel set of experiments using a more naturalistic word-level vocabulary. In case of LLMs, the distinction between a token-level and a word-level task may be less sharp than it appears. Kaplan et al. [2025] shows that LLMs form an “inner lexicon” where simple or common words are processed as cohesive semantic units. Given their results, this word-level task formulation can be considered analogous to the character-level task.
Experimental Setup
The setup for the word-level vocabulary mirrors the main experiments in Section 4.1 but replaces the character-based vocabulary with a vocabulary of 300+ unique English words. This allowed us to construct longer, more complex sequences with lengths up to 300 tokens. All other aspects of the methodology, few-shot prompting strategy, and task structures (e.g. UR vs. UL, UF vs. UB) remain identical. The exact vocabulary we used to generate the test strings in all of the experiments is as follows.
Character-level vocabulary
a, à, â, ā, ä, ą, b, c, ç, ć, ĉ, ċ, č, d, e, é,
ê, ë, ě, ē, ę, f, g, ĝ, ğ, h, ĥ, i, í, j, ĵ, k,
l, m, n, o, ó, ô, ö, õ, p, q, r, s, t, u, ú, ù,
û, ü, v, w, x, y, z,
A, Á, Â, Ã, Ä, Å, Æ, B, C, Ç, Ć, Ĉ, Ċ, Č, D, E,
È, É, Ē, Ė, F, G, H, I, Í, Ì, Î, Ï, Ī, Ĭ, Ĩ, J,
K, L, Ĺ, M, N, Ń, O, Ó, Ô, Ö, P, Q, R, S, T, U,
Ú, Ù, Û, Ü, V, W, X, Y, Z,
0, 1, 2, 3, 4, 5, 6, 7, 8, 9
Word-level vocabulary
apple, ant, arrow, anchor, artist, animal, angle, apricot, arch, armor, axis, avenue
ball, bat, book, bridge, bottle, bucket, bench, bread, bell, button, brush, branch
cat, car, cup, cloud, clock, candle, coin, chair, circle, crown, castle, cookie
dog, door, desk, drum, duck, doll, diamond, dish, dress, dream, drop, dust
egg, ear, eye, engine, elbow, earth, envelope, exit, echo, edge, event, energy
fish, fan, fork, flower, flag, feather, fire, frame, forest, farm, fruit, fence
goat, game, glass, glove, gate, garden, gift, grape, guitar, gold, gear, group
hat, hand, horse, house, hill, hammer, heart, honey, hook, horn, hug, hope
ice, iron, ink, island, idea, image, item, ivory, icon, input, issue, idol
jar, jam, jet, jewel, jungle, jacket, juice, job, joke, joy, judge, jump
kite, key, king, knee, kitchen, knife, kitten, knight, kick, kettle, kind, koala
lamp, leaf, lion, lock, ladder, lake, lemon, line, letter, lip, light, lunch
man, map, moon, milk, mouse, mirror, mountain, market, meal, music, magnet, match
net, nose, nest, name, nail, night, number, note, neck, nurse, noise, nation
owl, oil, oven, orange, ocean, order, orbit, open, option, owner, object, office
pen, pig, pot, plate, plane, pumpkin, pearl, park, path, piano, point, paper
queen, quill, quiz, quilt, quiet, quick, quote, quest, queue, quake, quart, quark
rat, ring, rain, river, rope, road, rose, rock, rule, room, root, radio
sun, sock, star, ship, shoe, stone, sugar, song, salt, sand, seed, snake
top, toy, tree, train, table, tube, tiger, tool, time, tent, team, towel
urn, use, unit, uncle, under, upper, uniform, union, urban, urge, ultra, usual
van, vase, veil, voice, valley, visit, value, vest, vote, view, vine, victory
wax, web, wall, wind, water, wheel, wave, wolf, wing, worm, word, wood
xylophone, xerox
yam, yard, yarn, yawn, year, yellow, yogurt, yolk, youth, yield, yeti, yoga
zoo, zebra, zone, zero, zip, zinc, zeal, zest, zigzag, zoom, zombie, zodiac
C.4 Additional results and plots for in-context prompting (Section 4.1)
C.4.1 Additional results: character-level vocabulary
Figure 10: In-context accuracy for Qwen2.5‑7B and Llama‑3-8B across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 11: In context accuracy for Qwen2.5‑7B-Instruct and Llama‑3-8B-Instruct across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 12: In context accuracy for Qwen2.5‑32B-Instruct and Llama‑3-70B-Instruct across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
As evident from all the figures here (smaller completion models - Figure 10), (smaller instruct models - Figure 11), (bigger instruct models - Figure 12), the Directional Bias persists across models sizes and types. While in the smaller models the gaps smaller for the retrieval tasks, nevertheless there is still a pronounced gap within the copying task variants.
Amongst the prompts there was high variability for all our retrieval setups, but close to none for the copying setups. Our bare template performed the worst in most retrieval setups, and the highest accuracy metrics were achieved when the instruction template was – MATH_RULE_EXPLAINED, with the few shot setting as SMALL with the presence of a separator. For copying, just the BARE setup was enough to get high accuracy numbers without the need to specify any instructions (especially for the forward copying cases). Providing instructions did improve the performance especially for the Qwen2.5-Instruct models.
We also carried out the same experiments with newer Qwen3 [Team, 2025] models with “thinking” variants to see if our findings still hold. The results from Figures 13, 14, 15, 16 show that even the model trained to explicitly generate a chain-of-thought for solving task still have the same Directional and Uniqueness biases.
Figure 13: In-context accuracy for Qwen3‑14B across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 14: In-context accuracy for Qwen3‑4B across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 15: In context accuracy for Qwen3‑4B-Thinking across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 16: In context accuracy for Qwen3‑14B-Thinking across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
C.4.2 Results: word-level vocabulary
We find that the core asymmetries persist even in the more naturalistic word-level setting (see Figures 17, 18, 19, 20, 21, 22, 23, 24). Although length generalization improves slightly for the larger models, the pretraining induced Directional Bias and the architecture-driven Uniqueness Bias remain clearly evident across all LLMs.
Figure 17: In context accuracy for Qwen2.5‑32B and Llama‑3-70B across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 18: In-context accuracy for Qwen2.5‑7B and Llama‑3-8B across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 19: In context accuracy for Qwen2.5‑7B-Instruct and Llama‑3-8B-Instruct across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 20: In context accuracy for Qwen2.5‑32B-Instruct and Llama‑3-70B-Instruct across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 21: In-context accuracy for Qwen3‑14B across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 22: In-context accuracy for Qwen3‑4B across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 23: In context accuracy for Qwen3‑4B-Thinking across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 24: In context accuracy for Qwen3‑14B-Thinking across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
C.5 Promping: Lorem Ipsum Copying (Section 4.2)
Dataset generation.
To stress‑test copying while keeping text realistic, we construct a 1,500‑example corpus of synthetic Lorem Ipsum paragraphs. Generation starts from one base paragraph produced by the lorem Python library and then applies light, probabilistic perturbations to increase the non-determinism of the texts.
  • Sentence count. Each sample contains exactly 45 sentences (average ~350 tokens), enforced by iterative expansion/shuffling until the target length is reached.
  • Word/discourse noise. With probability $0.3$ a random sentence is repeated up to four times; with probability $0.5$ random words inside a sentence are duplicated; with probability $1.0$ words within a sentence are shuffled (except the leading capital).
  • Token budget. Sequences are truncated to 500 BPE tokens to stay within context limits while preserving paragraph‑level coherence.
Prompt variants.
We identified successful prompts for copying naturalistic text as follows. We downloaded research papers put up on ArXiv in April 2025, and checked if the prompt results in perfect accuracy (not even a single mistake). Our motivation was that ArXiv text includes challenging text with equations, and, due to recency, should not have appeared in the LLM training data. More specifically, we segregated the sections of the papers into paragraphs, making sure the total length of the paragraph to be copied has close to 500 tokens (while ensuring that we do not end the paragraph in the middle of a sentence). We then pre-validated our prompts on a dataset containing $500$ samples each containing $500$ tokens, and even the smaller models we test – Llama3.1-8B and Qwen2.5-7B perfectly copied the given text (not a single mistake). We proceeded with our 3 minimal prompts, shown in Table 9, which achieved perfect copying accuracy. Each of our prompts ends with an unresolved <start> token, prompting the model to emit a verbatim repetition of the preceding paragraph.
Ratio of Ambiguous vs. Unambiguous Tokens. Table 7 reports, for each model, the average number of ambiguous (tokens with multiple possible continuations) and unambiguous (tokens with a single continuation) tokens per paragraph, aggregated across all random seeds. While different model families employ distinct tokenizers, so their absolute counts vary, yet the relative balance between ambiguous and unambiguous tokens is remarkably consistent. Each paragraph contains roughly 500 tokens, but repetition of sentences (to make the dataset appropriate for Non-unique copying) reduces the number of unique token types to only about 70. However, amongst those, both token types occur frequently in our data and the fact that copying “glitch” can be traced exclusively to unambiguous tokens is remarkable.
Table 7: Average per‐paragraph bigram counts by model
Model name Ambiguous tokens Unambiguous tokens Total tokens
Llama3_8B, Llama3_70B 31.97 39.65 71.62
Qwen2.5_7B, Qwen2.5_32B 32.59 42.14 74.72
Example of a Hallucinated Chain.
We provide here a truncated real sample (from the total 500 sequence length of the original paragraph) of an input sequence to be copied:
... neam que Non etincidunt dolorem tempora magnam.
and the corresponding output we get for the same input is:
... neam que Non etincidunt dolorem tempora magnam velit neque.
Non etincidunt dolorem tempora magnam.
As can be seen, a long chain of hallucination starts at the point tempora magnam. However, earlier in the truncated history, there were cases where magnam was followed by velit.
Our algorithm proceeds with this example as follows. We first tokenize both sequences to get the following lists of tokens in the input and the output. Here, the symbol $\dot{G}$ denotes a leading space for any token:
Input:  [‘Ġne’, ‘am’, ‘que’, ‘.’, ‘ĠNon’, ‘Ġet’, ‘inc’, ‘idunt’, ‘Ġdol’, ‘orem’,
‘Ġtemp’, ‘ora’, ‘Ġmagn’, ‘am’, ‘.‘],
Output: [‘Ġne’, ‘am’, ‘que’, ‘.’, ‘ĠNon’, ‘Ġet’, ‘inc’, ‘idunt’, ‘Ġdol’, ‘orem’,
         ‘Ġtemp’, ‘ora’, ‘Ġmagn’, ‘am’, ‘Ġvelit’, ‘Ġne’, ‘que’, ‘.’,     ‘ĠNon’,
         ‘Ġet’, ‘inc’, ‘idunt’, ‘Ġdol’, ‘orem’, ‘Ġtemp’, ‘ora’, ‘Ġmagn’, ‘am’, ‘.‘]
The alignment we obtain between input and output tokens is:
[(‘match’ 0, 0), (‘match’ 1, 1), (‘match’ 2, 2), (‘match’ 3, 3),
 (‘match’ 4, 4), (‘match’ 5, 5), (‘match’ 6, 6), (‘match’ 7, 7),
 (‘match’ 8, 8), (‘match’ 9, 9), (‘match’ 10, 10), (‘match’ 11, 11),
 (‘match’ 12, 12), (‘match’ 13, 13), (‘insert’, None, 14), (‘insert’, None, 15),
 (‘insert’, None, 16), (‘insert’, None, 17), (‘insert’, None, 18),
 (‘insert’, None, 19), (‘insert’, None, 20), (‘insert’, None, 21),
 (‘insert’, None, 22), (‘insert’, None, 23), (‘insert’, None, 24),
 (‘insert’, None, 25), (‘insert’, None, 26), (‘insert’, None, 27)]
After grouping the alignment by operation, we get:
[(‘match’ (0, 13)), (‘insert’, [14, 27))]
Finally, we analyze the transition index between these groups, which here is the index $13$. The token at that index is the word am, which appeared earlier in the context and was followed by other tokens and not a period (.). Since this transition index at am is ambiguous, we classify the source of this hallucination chain as ambiguous.
Table 8: Model Performance at Longer (2k/3k/5k) Input Lengths
Unambiguous Ambiguous
Model 2k 3k 5k 2k 3k 5k
llama3-70B-Instruct 0.99 0.99 0.99 0.96 0.94 0.76
llama3-8B-Instruct 0.96 0.98 0.99 0.89 0.69 0.44
qwen2.5-32B-Instruct 1.00 0.99 0.99 0.95 0.94 0.93
qwen2.5-7B-Instruct 0.99 0.99 0.99 0.75 0.43 0.19
Table 9: Prompt templates for copying Lorem Ipsum styled text. input stands for the test paragraph. All templates terminate immediately after the final <start> token, leaving the closing tags for the model to complete. The example paragraphs along with their copies are written in full in the templates for few-shot examples (omitted here for brevity)
Name Template body
echo‑tag RULE : Paragraphs are repeated **exactly** as is, whenever an Echo is seen.
Paragraph:
<start> Once upon a time … <end>
Echo:
<start> Once upon a time … <end>
Paragraph:
<start> input <end>
Echo:
<start>
verbatim‑end Paragraph:
<start> input <end>
Provide a verbatim copy of the paragraph above exactly as it appears, with no modifications. Use the tag <end> to mark the end of the response.
<start>
p‑tag‑the_end Please repeat the following paragraph exactly as it is, without any changes:
input
Put the repeated paragraph between <paragraph> and </paragraph> tags. Use the phrase "THE_END" to mark the end of the response.
<paragraph>
C.5.1 Additional results for longer lengths
We conducted additional experiments with the same Lorem-Ipsum setup described above for a single seed at even longer input lengths (up to 5000 tokens) and report the results in Table 8. We observe the same effects we see at lengths of 500, that is, a lower success rate for ambiguous bigrams than for unambiguous ones. These results on longer inputs align perfectly with and amplify our findings in Section 4.2.
C.6 Prompting: Git Commit Revert vs Cherry-Pick (Section 4.2)
Dataset generation.
[…]
All files are distributed in jsonl format with keys snippet, revert, and cherrypick.
Prompt variants.
We adopt exactly two natural‑language templates; one for the cherry‑pick orientation, one for the revert orientation. Each prompt describes the task in plain English, includes up to three few‑shot examples (hand‑crafted, $\ell=6$), shows the new history under a === NEW HISTORY === delimiter, and terminates with === ANSWER ===, where the model must emit the ordered commit list only. An example is given below.
Prompt template examples - Cherry Pick
Below is a part of a Git commit history (newest at the top, oldest at the bottom).
Task
List every commit line that must be cherry‑picked onto the release branch.
 Output them **from oldest to newest** (the order ‘git cherry‑pick‘ expects).
One complete commit line per output line.
Do **not** output anything else.
{{ few_shot_block }}
=== NEW HISTORY ===
{{ snippet }}
=== ANSWER ===
<start>
Prompt template examples - Revert
Below is part of a Git commit history (newest at the top, oldest at the bottom).
Task
List every commit line that must be reverted to roll the codebase back.
Output the lines **from newest to oldest**.
One complete commit line per output line.
Do **not** output anything else.
{{ few_shot_block }}
=== NEW HISTORY ===
{{ snippet }}
=== ANSWER ===
<start>
Because correct output demands global re‑ordering rather than token‑level copy, we found no benefit in adding separator toggles or mathematical re‑statements.
Model coverage.
Preliminary sweeps revealed that Qwen2.5 7B/32B models performed really poorly $\leq 10\%$ on this task, even at further shorter depths $\ell=5$, hence we exclude them from our analysis. Performance degrades further with longer histories, indicating a difficulty with list reversal rather than context window length.
Additional prompting details for Instruction-tuned model variants.
[…]
You are a very careful and precise assistant. You always follow the instructions
and solve tasks yourself. You never generate code.
With our initial experiments, we observed that unless we specify explicitly, instruction-tuned model variants generated code to solve both copying and retrieval tasks.
Appendix D Fine-Tuning Details (Section 4.3)
In general, the random seeds only affect the data shuffling and the positional offset during fine-tuning; the rest of the things remain unaffected.
D.1 Retrieval Tasks - UL/UR
Examples.
Some sample inputs with the query token (bolded) and their UL/UR answers
Input UL answer UR answer
ns0w6up9v8||u 6 p
qyw283zd9411w8||3 8 z
Dataset.
The training/test dataset statistics for the UL/UR tasks are as follows:
UL UR
# train/val/test samples 45k/5k/5k 45k/5k/5k
train/val sample lengths [min, max] [4, 100] [4, 100]
test sample lengths [min, max] [101, 200] [101, 200]
Hyperparameters.
The table below contains all the hyperparameters for fine-tuning models on the UL/UR dataset.
Hyperparameter Value
Model name GPT-2-XL [^12] https://huggingface.co/karpathy/gpt2_1558M_final4_hf
Tokenization Character-level
Maximum train sequence length 100
Maximum inference sequence length 200
# Epochs 30
Batch Size 64
Learning rate (AdamW) 1e-5
Weight decay (AdamW) 0.01
Warmup ratio 0.15
Random seeds 3, 71, 92, 435, 541, 591, 24050, 29214
Maximum gradient norm 1.0
D.2 Retrieval Tasks - NLFirst/NRFirst/NLLast/NRLast
Examples.
Some sample inputs with the query token (char after ‘||’) and their NLFirst/NRFirst/NLLast/NRLast targets.
Input NLFirst NRFirst NLLast NRLast
q5o0o8b6v5o3||o 5 0 5 3
c8r5r5r3r6r0||r 8 5 6 0
Dataset.
The training/test dataset statistics for the NLFirst/NRFirst/NLLast/NRLast tasks are as follows:
NLFirst/NRFirst/NLLast/NRLast
# train/val/test samples 100k/5k/4.7k
train/val sample lengths [min, max] [4, 100]
test sample lengths [min, max] [101, 200]
Hyperparameters.
The table below contains all the hyperparameters for fine-tuning models on the NLFirst/NRFirst/NLLast/NRLast dataset.
Hyperparameter Value
Model name GPT-2-XL
Tokenization Character-level
Maximum train sequence length 100
Maximum inference sequence length 200
# Epochs 15
Batch Size 64
Learning rate (AdamW) 1e-5
Weight decay (AdamW) 0.01
Warmup ratio 0.15
Random seeds 3, 47, 71, 100
Maximum gradient norm 1.0
D.3 Copying Tasks - UF/UB/NF/NB
Examples.
Some sample inputs and their UF/UB/NF/NB targets.
Input Config Target
Syb5DEHihO> UF Syb5DEHihO
Syb5DEHihO> UB OhiHED5byS
LnvTs1qgMt> UF LnvTs1qgMt
LnvTs1qgMt> UB tMgq1sTvnL
9975813713> NF 9975813713
9975813713> NB 3173185799
525671167> NF 525671167
525671167> NB 761176525
Dataset.
The training/test dataset statistics for the UF/UB/NF/NB tasks are as follows:
UF/UB/NF/NB
# train/val/test samples 50k/5k/5k
train/val sample lengths [min, max] [4, 100]
test sample lengths [min, max] [101, 200]
The train/val sample lengths are considered without the delimiter but the combined input/target length is 100, not just the input length.
Hyperparameters.
The table below contains all the hyperparameters for fine-tuning models on the UF/UB/NF/NB dataset.
Hyperparameter Value
Model name GPT-2-XL
Tokenization Character-level
Maximum train sequence length 100
Maximum inference sequence length 200
# Epochs 15
Batch Size 64
Learning rate (AdamW) 1e-5
Weight decay (AdamW) 0.01
Warmup ratio 0.15
Random seeds 3, 33, 71, 77, 91
Maximum gradient norm 1.0
D.4 Results
We provide the training loss curves for each fine-tuning task in Figure 25.
Figure 25: Averaged training loss curve for GPT-2-XL model across all our tasks. For all the tasks, we achieve 0 training loss. Yet, the generalization behavior varies in line with theoretical predictions (Figure 6).
Appendix E Further Details on Mechanistic Interpretability Analysis (Section 4.4)
Dataset.
We randomly generate 100 strings containing letters of the English alphabet in lowercase. To prompt base models to perform the tasks, we add 10 examples before the actual query for few-shot learning. We tokenize each symbol individually, adding a preceding white space to each letter to get a more natural tokenization. We add a BOS token in the beginning, separate few-shot examples with a dot, and separate the input and output of each example with a comma. Fine-tuned models use the same dataset, except that tokenization and the choice of separators are inherited from the fine-tuning setup, and they also have no BOS token and do not require few-shot examples.
Patching details.
To remove an induction head, we use path patching methodology, widely adopted in the literature [Hanna et al., 2023, Wang et al., 2022]. An induction head can be viewed as a path consisting of two edges connecting different positions inside an attention head: one edge connecting two adjacent positions in the input string, and one edge connecting the second of these with a position in the target string. We remove paths inside attention heads following Bakalova et al. [2025]. Removing all induction heads is thus done in two forward passes:
  • In all heads, replace the V activation of a token at position $P$ in the input string with zero whenever it is queried by the token in position $P+1$ in the input string. In the same forward pass, for each head, save K and V activations at position $P+1$ in the input string. Intuitively, this intervention prepares $K$ and $V$ activations at each position that carry no information about the preceding token.
  • In a second forward pass, in each head, replace the K activation of the token in position $P+1$ in the input string with the activation saved in the previous step whenever it is queried by a token in position $P$ in the target string. Intuitively, this intervention ensures that a head cannot attend to a position on the basis of information about the token immediately preceding the key position. That is precisely the behavior of induction heads [Olsson et al., 2022], i.e., induction heads are removed by this intervention.
When removing anti-induction heads, the first step is the same, but the second step is different: In each heads, we replace the V activation of the token in position $P+1$ in the input string with the activation saved in the first step whenever it is queried by a token in position $P+1$ in the target string. Intuitively, this intervention ensures that an attention head attending to a prior occurrence of the same token cannot retrieve the immediately preceding token.
The above description applies to Unique Forward Copy; in Unique Backward Copy, the position indices in the target string are appropriately reversed.
Layer localization.
Figure 26: Difference in sum of attention scores per layer between fine-tuned and base models. The higher the difference, the more amplified is this head in this layer during fine-tuning.
In fine-tuning, useful heads may migrate upward in depth (Figure 26). In fine-tuned models, induction (resp. anti-induction) strength peaks in the top-third layers for forward (resp. backward) copy. Pretrained models exhibit the same pattern but with induction heads situated later than anti-induction heads, mirroring their superior robustness.
Relation to Retrieval Heads.
Recent work by Wu et al. [2025] introduced the notion of a retrieval heads. The experiments to find these retrieval heads in Wu et al. [2025] focus primarily on zero-shot settings compared to the typical few shot settings for induction head studies [Elhage et al., 2021, Olsson et al., 2022, Song et al., 2025, Crosbie and Shutova, 2025] including our setup. These retrieval heads are found by looking at tokens being retrieved from Needle in a Haystack like test beds [Liu et al., 2024b]. There is a needle, or a short answer span that needs to be copied in the output given a query. These answer spans / needles consist of tokens that are sufficiently unique in the haystack (context) and thus cannot be figured out on the basis of semantics alone. A head is called a retrieval head if it pays maximal attention to tokens being copied from a needle more number of times than any other head. For example, if the number of tokens to be copied from a needle is 10, and the maximal retrieval score that was calculated came out to be 0.7, that would imply that the head in question paid the maximal attention to the 7 out of the 10 tokens while they were being copied. Thus even though Retrieval heads are highly connected to the way we calculate our induction and anti-induction circuits, the exact methodology is different. However retrieval heads should have a strong connection with the induction heads, when they perform forward copying, and the anti-induction heads when they perform reverse copying. The exact correlation however is left for future work.
Appendix F Compute Resources
We run all the prompting experiments using 4B/7B/8B parameter models on a single H100 GPU and 14B/32B/70B parameter models on 4xH100 GPUs with full precision without any quantization. Our fine-tuning experiments also use a single H100 GPU for all the experiments. All of our experiments should be reproducible with approximately 2500 H100 GPUh.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction assert that (1) fundamental length-generalization limits of small transformers may persist in large‐scale pretrained LLMs (2) pretrained models exhibit a notable induction-vs-anti-induction asymmetry; (3) this asymmetry can be traced to the relative strength of underlying attention circuits; and (4) targeted fine-tuning can rebalance these circuits and mitigate architectural limits, but does not entirely eliminate them. Through a mixture of synthetic as well as real world experiments in section 4.1, 4.2 we show point 2. Through our finetuning experiments in section 4.3 we justify point 4. We justify point 3 with experiments and discussion in section 4.4. Finally in totality the cumulation of all of our experiments and theoretical results justifies point 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations of our work in Section 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We base our theory on the extension of Huang et al. [2025]. In the main body of the paper we provide informal theorems and intuitions behind our ideas (Section 3). Detailed definitions, proofs, and remarks follow in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide technical details of our tasks, in-context and fine-tuning experiments in Section 4. Specifically, we address the setup of the copying tasks in subsection 4.2 and fine-tuning setup in subsection 4.3. Additional details on prompting and fine-tuning are provided in the Appendix C and D respectively.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We attach the code and data used all across our experiments in the supplementary section. We also have provided sufficient details all throughout the paper in each section to help the readers reproduce any result.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have provided exact hyperparameters in the appendix, along with all the relevant training and test details for all experiments considering training in the appendix in sections D, B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide exact error bars in our experiments whenever relevant. We use at least 3 seeds for each experiment of ours to confirm their validity.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer:[Yes]
  39. Justification: We provide details of the compute resources used for all our experiments in our Appendix in Section F.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: There is no deviation from the NeurIPS Code of Ethics guidelines mentioned. We conform to all the guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our paper as such connects theory of Transformers and contextualizes them for LLMs. Although we do not foresee any immediate negative positive societal impacts of our work. Long term, we hope that our insights that can be used to make better LLMs that when deployed in some critical domains can avoid the kind of errors we expose in LLMs
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A] .
  54. Justification: We do not foresee any misuse of the new benchmarks we release to expose some limitations in LLMs, and thus do not see the need to provide any safeguards related to the same.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the creators of models and code we use.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We contribute some new small benchmarks related to better understand limitations in LLMs, and attach the benchmarks, along with how to re-generate them in the supplementary material. We also attach all the code used to run our experiments in the supplementary section.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A] .
  69. Justification: We conduct no such experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A] .
  74. Justification: We conduct no such experiments.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Our paper is about studying LLMs, and hence all of our experiments involved running LLMs to understand their behavior. Additionally we used LLMs to assist in refining our writing throughout the paper. LLMs were not used in any other part of the paper, including our theoretical results, coming up with our experimental design, creating our datasets and writing code.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.